# XARP: A Human-First and AI Agent-Ready Extended Reality Toolkit

### v2

@@ Abstract
keywords:
Toolkit, XR, AI, Agents, Model Context ProtocolIntroduction
Extended reality (XR) systems increasingly integrate artificial intelligence (AI) to support task execution and skill acquisition in dynamic spatial environments. Recent AI-XR systems, for example, guide users through motor tasks with step-by-step instructions [47, 31] and provide personalized sports performance feedback [36, 34]. Yet designing such systems is challenging because current tools overburden developers with implementation details that divert attention from core design decisions of the AI-XR domain [25, 37]. Existing toolchains lack abstractions that bridge the heterogeneous demands of realtime interactivity on lightweight devices and large-scale inference on GPU clusters. This gap increases the risk of ad hoc integrations across rendering engines and AI services, making systems fragile to the rapid turnover of SDKs and frameworks in this domain. As a result, implementation and deployment become laborious and error-prone, draining resources that could instead address essential design challenges of AI-XR systems [5, 6], such as user interaction and contextual understanding in unseen environments [54, 10]. Higher-level abstractions that unify access to XR and AI capabilities could reduce this burden, enabling developers to redirect effort toward rigorous evaluation and innovative AI-XR designs.
@@
Consider a fitness application built with XARP. Developers can call high-level XR functions to guide users through calibration, assess available training space, and track skill level through body pose estimation. At runtime, AI agents can reuse these functions to adapt exercises dynamically based on user performance, environmental changes, or new user preferences. Through MCP integration, third-party providers could plug in AI coaches with distinct training styles that seamlessly extend the application. Unlike existing solutions [35, 54, 20, 31], XARP’s unified abstraction simultaneously reduces the implementation burden on developers and enables them to delegate design decisions to AI agents at runtime, trading specification effort for adaptive, dynamic behavior. XARP is the first system to expose XR capabilities through MCP, simplifying AI-XR integration with an open standard.
To demonstrate the feasibility of XARP and its ability to support diverse AI-XR applications, we conducted three case studies: AI-guided fencing, drone assistance, and room layout design. Design reflections after each case study helped align XARP with how developers actually build AI-XR systems. These studies surfaced recurring challenges and common development roles, motivating features such as persistent interaction logs visualized in a chat format, rapid prototyping via a web client, and separation of concerns in team workflows through clear boundaries between front- and back-end experts. We evaluated adoption potential through a video walkthrough study with 24 AI and XR developers. Participants watched step-by-step demonstrations of seven tasks interleaved with comprehension checks. The tasks covered the toolkit setup, extension, and all three modes of use. Afterward, they completed a survey based on established toolkit evaluation goals [28] and the UTAUT technology adoption scale [48]. In open-ended responses, participants described when they would or would not use XARP and posed questions that remained after the walkthrough.
@@ 1 Toolkit Architecture
1.1 Backend and Frontend Separation
XARP adopts a clear architectural separation between backend application logic and frontend XR clients, enabling flexible deployment and scalable team collaboration. This separation allows developers to iterate on backend logic without rebuilding or redeploying client applications, accelerating experimentation and reducing development overhead. The architecture also supports modular, role-specific workflows aligned with Conway’s Law [13], which suggests that system designs tend to mirror the communication structures of the organizations that produce them. Teams can divide responsibilities across XR specialists, Python developers, and AI engineers, each working within their domain without conflicting dependencies. This architecture design enables multi-platform deployment, where any XR client that implements the required set of primitive tools can connect to the same backend application. This architecture also supports multi-client experiences, where different clients connect to the same backend app and share data with a session key. This is useful to support copresence, with multiple users connected from different devices. Another advantage of this design is the cross-device synergy, e.g., pairing a VR headset with a mobile device to enable efficient text entry or multiview video.
1.2 High-level XR Functions
@@
1.3 AI Agent Integration
XARP enables AI agents to engage directly in XR experiences beyond coding assistance or asset generation, as active participants during runtime. Agents gain access to the same high-level XR functions available to human developers, exposed as callable tools through a shared abstraction layer. To support generalization across platforms, XARP includes a capability discovery mechanism where each connected XR client declares its supported tools to the server, allowing AI agents to adapt their behavior dynamically based on different client capabilities. This mechanism can potentially enable the same application logic to operate across heterogeneous devices, from high-end XR headsets to mobile AR platforms. By supporting runtime behavior generation, XARP reduces the need for complete system specification at design time with AI-driven adaptiveness. The Model Context Protocol (MCP) [2] is an open standard for connecting AI agents to external resources, executable tools, and reusable prompts. While MCP introduces challenges such as stateful session management and external system dependencies, it represents an important step toward standardized AI integration. XARP is the first toolkit to expose XR functionalities via MCP, allowing third-party agents to interact seamlessly with users in XR. This capability expands the design space for AI-XR applications and contributes to the convergence of these ecosystems.
1.4 Implementation Details
@@
2 Related Work
Drawing from seminal work in software engineering, we divide the challenges in AI-XR development into essential and accidental complexity [25, 37]. Essential complexity refers to the intrinsic difficulty of conceptualizing a solution, while accidental complexity arises from translating that solution into a concrete implementation using the available tools and processes [25, 37]. Prior work has tackled both forms of complexity through toolkits, asset generation pipelines, and adaptive systems.
2.1 Reducing Accidental Complexity in AI-XR
To reduce accidental complexity, vendors integrate their XR platforms with AI services, such as Android XR with Gemini and Snap OS with Lens Studio AI. While these solutions lower development friction within proprietary ecosystems, they introduce vendor lock-in. Unity Sentis mitigates this by relying on the ONNX open standard to enable on-device AI, though execution remains bound to the Unity engine. In a different strategy, AI has been used to reduce accidental complexity by automating code generation [49, 24, 40] and accelerating asset creation [52, 26]. But these approaches address only a narrow part of the entire development pipeline and require manual effort to integrate output into XR applications. Specialized middleware such as Estuary [35], CUIfy [8], and NVIDIA ACE further reduce authoring effort costs for conversational embodied avatars, but are not designed for general-purpose AI-XR development.
@@ 2.1 Reducing Accidental Complexity in AI-XR
2.2 Managing Essential Complexity in AI-XR
Efforts to manage essential complexity have followed two main strategies: design toolkits and adaptive systems. Design toolkits support early-stage analysis and prototyping [10, 55], but alone do not produce deployable solutions. Adaptive systems shift decisions from specification time to runtime, allowing systems to respond to the concrete context of use[18, 33]. However, most adaptive XR systems remain limited to spatial layout and interface adjustments. More recent systems aim to bridge the gap between code, assets, and deployable XR applications, enabling runtime generation of end-user experiences [20, 14, 31]. Although these approaches can manage essential complexity by deferring design decisions to runtime, they are typically implemented as standalone systems tailored to specific use cases rather than general-purpose toolkits. An exception is XaiR [54], which offers a pathway for building new adaptive systems but does not yet provide a high-level abstraction usable by both developers and AI agents.
@@ 2.2 Managing Essential Complexity in AI-XR
3 Design Patterns from Literature
Inspired by prior work on design pattern elicitation in HCI and software engineering [4, 43, 50], we derived recurring interaction patterns from a review of recent AI–XR systems. Our process involved: (1) selecting a representative set of published systems across academic and industry sources; (2) identifying common functional features and user-agent interactions; (3) abstracting each feature into a design pattern described in terms of context, problem, and solution; and (4) iteratively refining the patterns.
@@ 3 Design Patterns from Literature
AI-XR systems must reliably capture user input that can be processed in real time by the application or delegated to an AI agent. Across prior work, three modalities recur most frequently: hand and gesture-based interaction [41, 23, 3, 44, 56, 22, 29, 16, 15, 1], eye and head gaze [11, 56, 22, 51, 12, 29, 7, 30], and speech input [3, 31, 47, 1, 30, 23]. Implementing these modalities typically requires juggling platform-specific SDKs and device APIs, creating fragile pipelines that complicate integration with AI models.
@@
Most intelligent XR systems aim to support users beyond executing explicit commands, and a key enabler of this functionality is providing AI models or agents with additional contextual information that can be used to predict user needs and offer proactive assistance. Common sources of such context include user position in the environment [33, 23, 11, 3, 47], RGB frames or continuous video streams from the device camera [12, 3, 33, 41, 11], and depth information aligned with those frames [11, 33, 47, 44, 3, 56]. Implementing these inputs often requires stitching together fragmented APIs and sensor pipelines, which increases engineering complexity and makes it difficult for AI agents to access context in a consistent format.
@@
Spatially aware and intelligent XR applications use contextual information not only to capture raw data but also to process it into higher-level representations that enable interactivity or support user goals. For example, systems can recognize objects in the environment [16] and infer their functionality, supporting tasks such as room layout rearrangement [1] and retrieval of information about shared objects in the space [53]. Other forms of context processing observed in prior work include measuring and modeling the 3D environment [29, 12, 51, 15, 44, 47, 3], as well as generating new 3D objects based on user needs and application goals [15, 1]. While these capabilities expand the scope of AI-XR interaction, they often require combining multiple perception pipelines and inference services, creating complex dependencies that are difficult for developers to implement and for AI agents to reuse.
@@
A fundamental component of interactive XR and AI systems is producing feedback that responds to user queries or commands, both to confirm whether actions were successfully executed and to convey additional information when needed. Prior work shows that such feedback spans multiple modalities. For non-spatial queries, responses are often delivered through audio [31, 47, 1, 7, 15, 16, 30, 29, 11] or text [11, 47, 22, 44, 3, 7, 16, 31]. When the feedback relates directly to the spatial environment, 3D annotations such as arrows, highlights, or bounding boxes are more effective at conveying system state and guiding user action [31, 47, 1, 7, 29]. Implementing these feedback channels typically requires a mix-and-match across separate XR toolkits and UI frameworks, complicating development and limiting how AI agents can dynamically tailor feedback to the situation.
@@
4 Case Studies
We conducted three in-house case studies using XARP to probe toolkit limitations and examine the AI-XR development process. These scenarios revealed recurring challenges, common developer workflows, and role divisions, which guided subsequent system refinements. The fencing coach case study emphasized instruction and feedback on the user’s motor actions. The drone assistance case study emphasized interaction with an embodied agent that paves the way toward robotic embodiments. The room layout case study emphasized virtual augmentations on physical space. Together, these case studies span essential elements of AI-XR design, i.e., the user, AI agent, and environment [10].
4.1 Fencing Coach
@@
Challenges and insights. Implementing the feedback system required capturing full-body pose data using sensors beyond the capabilities of standard XR headsets [32]. The integration of external RGB video for pose tracking and the calculation of joint angle features revealed two core development challenges: (1) the need for flexible sensor integration in XR environments, and (2) the importance of comprehensive data logging to support both AI grounding and developer troubleshooting. In particular, enabling AI agents to generate grounded feedback using Retrieval-Augmented Generation (RAG) workflows demanded persistent access to temporally indexed interaction traces.
@@ 4.2 Drone Assistant
Throughout three design iterations, end-users expected agents to understand gestures, speech, changes in the physical surroundings, and to provide visual feedforward of planned paths. To generate realistic flight behavior, we relied on AirSim [45]. Integrating this third-party component was a key challenge, as it uses Unreal as an underlying physics simulator, which brings a cascade of dependencies into the development workflow.
@@ 4.3 Room Layout
5 Walkthrough Study
Our goal in this study was to evaluate the perceived utility and adoption potential of XARP. Rather than measuring implementation-specific usability or performance metrics, which could obscure the evaluation of our underlying design strategy [21], we focused on how developers understand and assess the benefits of XARP for the development of future AI-XR systems. To this end, we conducted video walkthrough demonstrations [42, 17, 28] with 24 potential users. This method mirrors how developers often engage with new tools by watching tutorials before deciding whether to adopt them. It also echoes a long-standing tradition of using demonstrations to convey system potential, as seen in Engelbart’s 1968 “Mother of All Demos.” After the walkthrough, participants completed a survey combining a standardized technology adoption scale [48] with a custom questionnaire inspired by common goals of HCI toolkits [28] to assess XARP’s perceived value and potential for adoption in their own workflows.
5.1 Participants

### v3

@@ Abstract.
1. Introduction
Extended reality (XR) systems increasingly integrate artificial intelligence (AI) to support task execution and skill acquisition in dynamic spatial environments. Recent AI-XR systems, for example, guide users through motor tasks with step-by-step instructions (Srinidhi et al., 2024; Li et al., 2025b) and provide personalized sports performance feedback (Mandic et al., 2023; Liao et al., 2025). Yet designing such systems is challenging because current tools overburden developers with implementation details that divert attention from core design decisions of the AI-XR domain (Jr., 1987; Moseley and Marks, 2006). Existing toolchains lack abstractions that bridge the heterogeneous demands of realtime interactivity on lightweight devices and large-scale inference on GPU clusters. This gap increases the risk of ad hoc integrations across rendering engines and AI services, making systems fragile to the rapid turnover of SDKs and frameworks in this domain. As a result, implementation and deployment become laborious and error-prone, draining resources that could instead address essential design challenges of AI-XR systems (Börsting et al., 2022; Bose and Brown, 2024), such as user interaction and contextual understanding in unseen environments (Xu et al., 2023; Caetano et al., 2025). Higher-level abstractions that unify access to XR and AI capabilities could reduce this burden, enabling developers to redirect effort toward rigorous evaluation and innovative AI-XR designs.
@@ 1. Introduction
Consider a fitness application built with XARP. Developers can call high-level XR functions to guide users through calibration, assess available training space, and track skill level through body pose estimation. At runtime, AI agents can reuse these functions to adapt exercises dynamically based on user performance, environmental changes, or new user preferences. Through MCP integration, third-party providers could plug in AI coaches with distinct training styles that seamlessly extend the application. Unlike existing solutions (Lin et al., 2024; Xu et al., 2023; Giunchi et al., 2024; Li et al., 2025b), XARP’s unified abstraction simultaneously reduces the implementation burden on developers and enables them to delegate design decisions to AI agents at runtime, trading specification effort for adaptive, dynamic behavior. XARP is the first system to expose XR capabilities through MCP, simplifying AI-XR integration with an open standard.
To demonstrate the feasibility of XARP and its ability to support diverse AI-XR applications, we conducted three case studies: AI-guided fencing, drone assistance, and room layout design. Design reflections after each case study helped align XARP with how developers actually build AI-XR systems. These studies surfaced recurring challenges and common development roles, motivating features such as persistent interaction logs visualized in a chat format, rapid prototyping via a web client, and separation of concerns in team workflows through clear boundaries between front- and back-end experts. We evaluated adoption potential through a video walkthrough study with 24 AI and XR developers. Participants watched step-by-step demonstrations of seven tasks interleaved with comprehension checks. The tasks covered the toolkit setup, extension, and all three modes of use. Afterward, they completed a survey based on established toolkit evaluation goals (Ledo et al., 2018) and the UTAUT technology adoption scale (Venkatesh et al., 2003). In open-ended responses, participants described when they would or would not use XARP and posed questions that remained after the walkthrough.
@@ 2. Toolkit Architecture
2.1. Backend and Frontend Separation
XARP adopts a clear architectural separation between backend application logic and frontend XR clients, enabling flexible deployment and scalable team collaboration. This separation allows developers to iterate on backend logic without rebuilding or redeploying client applications, accelerating experimentation and reducing development overhead. The architecture also supports modular, role-specific workflows aligned with Conway’s Law (Conway, 1968), which suggests that system designs tend to mirror the communication structures of the organizations that produce them. Teams can divide responsibilities across XR specialists, Python developers, and AI engineers, each working within their domain without conflicting dependencies. This architecture design enables multi-platform deployment, where any XR client that implements the required set of primitive tools can connect to the same backend application. This architecture also supports multi-client experiences, where different clients connect to the same backend app and share data with a session key. This is useful to support copresence, with multiple users connected from different devices. Another advantage of this design is the cross-device synergy, e.g., pairing a VR headset with a mobile device to enable efficient text entry or multiview video.
2.2. High-level XR Functions
@@
2.3. AI Agent Integration
XARP enables AI agents to engage directly in XR experiences beyond coding assistance or asset generation, as active participants during runtime. Agents gain access to the same high-level XR functions available to human developers, exposed as callable tools through a shared abstraction layer. To support generalization across platforms, XARP includes a capability discovery mechanism where each connected XR client declares its supported tools to the server, allowing AI agents to adapt their behavior dynamically based on different client capabilities. This mechanism can potentially enable the same application logic to operate across heterogeneous devices, from high-end XR headsets to mobile AR platforms. By supporting runtime behavior generation, XARP reduces the need for complete system specification at design time with AI-driven adaptiveness. The Model Context Protocol (MCP) (Anthropic, 2024) is an open standard for connecting AI agents to external resources, executable tools, and reusable prompts. While MCP introduces challenges such as stateful session management and external system dependencies, it represents an important step toward standardized AI integration. XARP is the first toolkit to expose XR functionalities via MCP, allowing third-party agents to interact seamlessly with users in XR. This capability expands the design space for AI-XR applications and contributes to the convergence of these ecosystems.
2.4. Implementation Details
@@
3. Related Work
Drawing from seminal work in software engineering, we divide the challenges in AI-XR development into essential and accidental complexity (Jr., 1987; Moseley and Marks, 2006). Essential complexity refers to the intrinsic difficulty of conceptualizing a solution, while accidental complexity arises from translating that solution into a concrete implementation using the available tools and processes (Jr., 1987; Moseley and Marks, 2006). Prior work has tackled both forms of complexity through toolkits, asset generation pipelines, and adaptive systems.
3.1. Reducing Accidental Complexity in AI-XR
To reduce accidental complexity, vendors integrate their XR platforms with AI services, such as Android XR with Gemini and Snap OS with Lens Studio AI. While these solutions lower development friction within proprietary ecosystems, they introduce vendor lock-in. Unity Sentis mitigates this by relying on the ONNX open standard to enable on-device AI, though execution remains bound to the Unity engine. In a different strategy, AI has been used to reduce accidental complexity by automating code generation (Wermelinger, 2023; Jimenez et al., 2024; Nascimento et al., 2025) and accelerating asset creation (Xiang et al., 2025; Jun and Nichol, 2023). But these approaches address only a narrow part of the entire development pipeline and require manual effort to integrate output into XR applications. Specialized middleware such as Estuary (Lin et al., 2024), CUIfy (Buldu et al., 2025), and NVIDIA ACE further reduce authoring effort costs for conversational embodied avatars, but are not designed for general-purpose AI-XR development.
@@ 3.1. Reducing Accidental Complexity in AI-XR
3.2. Managing Essential Complexity in AI-XR
Efforts to manage essential complexity have followed two main strategies: design toolkits and adaptive systems. Design toolkits support early-stage analysis and prototyping (Caetano et al., 2025; Yildirim et al., 2023), but alone do not produce deployable solutions. Adaptive systems shift decisions from specification time to runtime, allowing systems to respond to the concrete context of use(Evangelista Belo et al., 2022; Li et al., 2024). However, most adaptive XR systems remain limited to spatial layout and interface adjustments. More recent systems aim to bridge the gap between code, assets, and deployable XR applications, enabling runtime generation of end-user experiences (Giunchi et al., 2024; De La Torre et al., 2024a; Li et al., 2025b). Although these approaches can manage essential complexity by deferring design decisions to runtime, they are typically implemented as standalone systems tailored to specific use cases rather than general-purpose toolkits. An exception is XaiR (Xu et al., 2023), which offers a pathway for building new adaptive systems but does not yet provide a high-level abstraction usable by both developers and AI agents.
@@ 3.2. Managing Essential Complexity in AI-XR
4. Design Patterns from Literature
Inspired by prior work on design pattern elicitation in HCI and software engineering (Borchers, 2000; Retalis et al., 2006; Winters and Mor, 2009), we derived recurring interaction patterns from a review of recent AI–XR systems. Our process involved: (1) selecting a representative set of published systems across academic and industry sources; (2) identifying common functional features and user-agent interactions; (3) abstracting each feature into a design pattern described in terms of context, problem, and solution; and (4) iteratively refining the patterns.
@@ 4. Design Patterns from Literature
AI-XR systems must reliably capture user input that can be processed in real time by the application or delegated to an AI agent. Across prior work, three modalities recur most frequently: hand and gesture-based interaction (Numan et al., 2025; Hu et al., 2025; Bohus et al., 2024; Rheault et al., 2024; Yoo et al., 2023; Hashky et al., 2024; Lee et al., 2024a; Dogan et al., 2024; De La Torre et al., 2024b; Aghel Manesh et al., 2024), eye and head gaze (Cai et al., 2025; Yoo et al., 2023; Hashky et al., 2024; Wu et al., 2024; Castelo et al., 2023; Lee et al., 2024a; Bovo et al., 2025; Lee et al., 2024b), and speech input (Bohus et al., 2024; Li et al., 2025b; Srinidhi et al., 2024; Aghel Manesh et al., 2024; Lee et al., 2024b; Hu et al., 2025). Implementing these modalities typically requires juggling platform-specific SDKs and device APIs, creating fragile pipelines that complicate integration with AI models.
@@
Most intelligent XR systems aim to support users beyond executing explicit commands, and a key enabler of this functionality is providing AI models or agents with additional contextual information that can be used to predict user needs and offer proactive assistance. Common sources of such context include user position in the environment (Li et al., 2024; Hu et al., 2025; Cai et al., 2025; Bohus et al., 2024; Srinidhi et al., 2024), RGB frames or continuous video streams from the device camera (Castelo et al., 2023; Bohus et al., 2024; Li et al., 2024; Numan et al., 2025; Cai et al., 2025), and depth information aligned with those frames (Cai et al., 2025; Li et al., 2024; Srinidhi et al., 2024; Rheault et al., 2024; Bohus et al., 2024; Yoo et al., 2023). Implementing these inputs often requires stitching together fragmented APIs and sensor pipelines, which increases engineering complexity and makes it difficult for AI agents to access context in a consistent format.
@@
Spatially aware and intelligent XR applications use contextual information not only to capture raw data but also to process it into higher-level representations that enable interactivity or support user goals. For example, systems can recognize objects in the environment (Dogan et al., 2024) and infer their functionality, supporting tasks such as room layout rearrangement (Aghel Manesh et al., 2024) and retrieval of information about shared objects in the space (Xu et al., 2024). Other forms of context processing observed in prior work include measuring and modeling the 3D environment (Lee et al., 2024a; Castelo et al., 2023; Wu et al., 2024; De La Torre et al., 2024b; Rheault et al., 2024; Srinidhi et al., 2024; Bohus et al., 2024), as well as generating new 3D objects based on user needs and application goals (De La Torre et al., 2024b; Aghel Manesh et al., 2024). While these capabilities expand the scope of AI-XR interaction, they often require combining multiple perception pipelines and inference services, creating complex dependencies that are difficult for developers to implement and for AI agents to reuse.
@@
A fundamental component of interactive XR and AI systems is producing feedback that responds to user queries or commands, both to confirm whether actions were successfully executed and to convey additional information when needed. Prior work shows that such feedback spans multiple modalities. For non-spatial queries, responses are often delivered through audio (Li et al., 2025b; Srinidhi et al., 2024; Aghel Manesh et al., 2024; Bovo et al., 2025; De La Torre et al., 2024b; Dogan et al., 2024; Lee et al., 2024b, a; Cai et al., 2025) or text (Cai et al., 2025; Srinidhi et al., 2024; Hashky et al., 2024; Rheault et al., 2024; Bohus et al., 2024; Bovo et al., 2025; Dogan et al., 2024; Li et al., 2025b). When the feedback relates directly to the spatial environment, 3D annotations such as arrows, highlights, or bounding boxes are more effective at conveying system state and guiding user action (Li et al., 2025b; Srinidhi et al., 2024; Aghel Manesh et al., 2024; Bovo et al., 2025; Lee et al., 2024a). Implementing these feedback channels typically requires a mix-and-match across separate XR toolkits and UI frameworks, complicating development and limiting how AI agents can dynamically tailor feedback to the situation.
@@
5. Case Studies
We conducted three in-house case studies using XARP to probe toolkit limitations and examine the AI-XR development process. These scenarios revealed recurring challenges, common developer workflows, and role divisions, which guided subsequent system refinements. The fencing coach case study emphasized instruction and feedback on the user’s motor actions. The drone assistance case study emphasized interaction with an embodied agent that paves the way toward robotic embodiments. The room layout case study emphasized virtual augmentations on physical space. Together, these case studies span essential elements of AI-XR design, i.e., the user, AI agent, and environment (Caetano et al., 2025).
5.1. Fencing Coach
@@
Challenges and insights. Implementing the feedback system required capturing full-body pose data using sensors beyond the capabilities of standard XR headsets (Li et al., 2025a). The integration of external RGB video for pose tracking and the calculation of joint angle features revealed two core development challenges: (1) the need for flexible sensor integration in XR environments, and (2) the importance of comprehensive data logging to support both AI grounding and developer troubleshooting. In particular, enabling AI agents to generate grounded feedback using Retrieval-Augmented Generation (RAG) workflows demanded persistent access to temporally indexed interaction traces.
@@ 5.2. Drone Assistant
Throughout three design iterations, end-users expected agents to understand gestures, speech, changes in the physical surroundings, and to provide visual feedforward of planned paths. To generate realistic flight behavior, we relied on AirSim (Shah et al., 2018). Integrating this third-party component was a key challenge, as it uses Unreal as an underlying physics simulator, which brings a cascade of dependencies into the development workflow.
@@ 5.3. Room Layout
6. Walkthrough Study
Our goal in this study was to evaluate the perceived utility and adoption potential of XARP. Rather than measuring implementation-specific usability or performance metrics, which could obscure the evaluation of our underlying design strategy (Greenberg and Buxton, 2008), we focused on how developers understand and assess the benefits of XARP for the development of future AI-XR systems. To this end, we conducted video walkthrough demonstrations (Ramakers et al., 2016; Evangelista Belo et al., 2021; Ledo et al., 2018) with 24 potential users. This method mirrors how developers often engage with new tools by watching tutorials before deciding whether to adopt them. It also echoes a long-standing tradition of using demonstrations to convey system potential, as seen in Engelbart’s 1968 “Mother of All Demos.” After the walkthrough, participants completed a survey combining a standardized technology adoption scale (Venkatesh et al., 2003) with a custom questionnaire inspired by common goals of HCI toolkits (Ledo et al., 2018) to assess XARP’s perceived value and potential for adoption in their own workflows.
6.1. Participants